# CodeGRAG: Extracting Composed Syntax Graphs for Retrieval Augmented Cross-Lingual Code Generation

### v1

@@ 4.4 Impacts of the Components of the Graphical View (RQ5)
In this section, we adjust the inputs of the graph components to the LLMs as the external knowledge. Concretely, we study the information contained in node names, edge names, and the topological structure. The results are presented in Table 3.
@@
4.5 Discussion of the Compatibility (RQ6)
@@
In addition, we give the ratio of generated results that can pass the graphical views extraction process, which is denoted by Extraction Rate. The Pass@1 and the Extraction Rate of the generated results passing the graphical extraction process are given in Table 4 and Table 5.
Table 4: The extraction rate of the generated results passing the graphical extraction process (C++).
@@
Code-RAG 62.20 92.07
CodeGRAG 64.02 92.68
Table 5: The extraction rate of the generated results passing the graphical extraction process (Python).
Generated Codes Pass@1 Extraction Rate
Code-RAG 71.95 91.46
CodeGRAG 77.44 96.95
From the results, we could see that the extraction rates are high for codes to pass the graphical views extraction process, even under the situation where the Pass@1 ratios of the generated results are low. This indicates that the application range of the proposed method is wide. In addition, as the code RAG also offers performance gains, one could use multiple views, including raw code blocks and graphical representations, as the retrieval knowledge base.
@@ 5 Conclusion
Despite the expanding role of LLMs in code generation, there are inherent challenges pertaining to their understanding of code syntax and their multi-lingual code generation capabilities. This paper introduces the Syntax Graph Retrieval Augmented Code Generation (CodeGRAG) to enhance LLMs for single round and cross-lingual code generation. CodeGRAG extracts and summarizes data flow and control flow information from codes, effectively bridging the gap between programming language and natural language. By integrating external structural knowledge, CodeGRAG enhances LLMs’ comprehension of code syntax and empowers them to generate complex and multi-lingual code with improved accuracy and fluency. The method’s demonstrated performance gains signal its potential to significantly elevate LLMs’ code generation abilities, including facilitating cross-lingual programming language generation such as C++ to Python. Overall, the findings warrant optimism regarding the transformative impact of CodeGRAG in advancing LLMs’ proficiency in code generation across various programming languages.
Limitations
In this paper, we propose a tuning-free method that can offer enhanced single round and cross-lingual code generation. Despite the efficiency and effectiveness, there are also limitations within this work. Firstly, dependency on the quality of the external knowledge base could be a potential concern. The quality of the external knowledge base could be improved with regular expression extraction on the noisy texts and codes. Secondly, the potential of combining the RAG with finetuning for improved generation is not explored in this work. Finetuning the LLMs with the assistance of CodeGRAG could be a future exploration.
Appendix A Example of the inserted graphical view
An illustration of the inserted graphical view is given below.
{mdframed}
Graph(
num_nodes=’node’: 24,
num_edges=(’node’, ’-0’, ’node’): 1, (’node’, ’-1’, ’node’): 1, (’node’, ’ArraySubscriptExpredge0’, ’node’): 1, (’node’, ’ArraySubscriptExpredge1’, ’node’): 1, (’node’, ’CXXOperatorCallExpredge1’, ’node’): 1, (’node’, ’CXXOperatorCallExpredge2’, ’node’): 2, (’node’, ’ImplicitCastExpredge0’, ’node’): 1, (’node’, ’UserDefineFun’, ’node’): 1, (’node’, ’falseNext’, ’node’): 1, (’node’, ’next’, ’node’): 5, (’node’, ’read’, ’node’): 10, (’node’, ’trueNext’, ’node’): 1, (’node’, ’write’, ’node’): 9,
metagraph=[(’node’, ’node’, ’-0’), (’node’, ’node’, ’-1’), (’node’, ’node’, ’ArraySubscriptExpredge0’), (’node’, ’node’, ’ArraySubscriptExpredge1’), (’node’, ’node’, ’CXXOperatorCallExpredge1’), (’node’, ’node’, ’CXXOperatorCallExpredge2’), (’node’, ’node’, ’ImplicitCastExpredge0’), (’node’, ’node’, ’UserDefineFun’), (’node’, ’node’, ’falseNext’), (’node’, ’node’, ’next’), (’node’, ’node’, ’read’), (’node’, ’node’, ’trueNext’), (’node’, ’node’, ’write’)]

### v2

@@ 3.4 Impacts of the Components of the Graphical View (RQ5)
In this section, we adjust the inputs of the graphical components to the LLMs. Concretely, we study the information contained in node names, edge names, and the topological structure. The results are presented in Table 4.
@@
3.5 Compatibility Discussion of the Graphical Views(RQ6)
@@
In addition, we give the ratio of generated results that can pass the graphical views extraction process, which is denoted by Extraction Rate. The Pass@1 and the Extraction Rate of the generated results passing the graphical extraction process are given in Table 5.
Table 5: The extraction rate of the generated results passing the graphical extraction process.
@@
(C++) Code-RAG 62.20 92.07
(C++) CodeGRAG 64.02 92.68
(Python) Code-RAG 71.95 91.46
(Python) CodeGRAG 77.44 96.95
From the results, we could see that the extraction rates are high for codes to pass the graphical views extraction process, even under the situation where the Pass@1 ratios of the generated results are low. This indicates that the application range of the proposed method is wide. In addition, as the code RAG also offers performance gains, one could use multiple views as the retrieval knowledge.
4 Related Work
Code Search. The code search methods can be summarized into three folds. Early methods utilizes sparse search to match the query and codes (Hill et al., 2011; Yang & Huang, 2017), which suffers from mismatched vocabulary due to the gap between natural language and codes. Neural methods (Cambronero et al., 2019; Gu et al., 2021) then focus on mapping the query and codes into a joint representation space for more accurate retrieval. With the success of pretrained language models, many methods propose to use pretraining tasks to improve the code understanding abilities and align different language spaces. For example, CodeBERT (Feng et al., 2020) is pretrained on NL-PL pairs of 6 programming languages with the masked language modeling and replaced token detection task. CodeT5 (Wang et al., 2021) supports both code-related understanding and generation tasks through bimodal dual generation. UniXcoder (Guo et al., 2022) integrates the aforementioned pretraining tasks, which is a unified cross-modal pre-trained model. As retrieval augmented generation (RAG) shows its significance in promoting the quality of LLMs generation, works in code RAG start to accumulate. Nashid et al. (2023); Lu et al. (2022) utilize the code blocks as the retrieved knowledge to inform the LLMs with similar code blocks for enhancement. Zhou et al. (2022) uses the programming related document to serve as the retrieval content, injecting auxiliary external programming knowledge into the LLMs generation.
@@ 5 Conclusion
Despite the expanding role of LLMs in code generation, there are inherent challenges pertaining to their understanding of code syntax and their multi-lingual code generation capabilities. This paper introduces the Syntax Graph Retrieval Augmented Code Generation (CodeGRAG) to enhance LLMs for single round and cross-lingual code generation. CodeGRAG extracts and summarizes data flow and control flow information from codes, effectively bridging the gap between programming language and natural language. By integrating external structural knowledge, CodeGRAG enhances LLMs’ comprehension of code syntax and empowers them to generate complex and multi-lingual code with improved accuracy and fluency.
Author Contributions
If you’d like to, you may include a section for author contributions as is done in many journals. This is optional and at the discretion of the authors.
Acknowledgments
Use unnumbered third level headings for the acknowledgments. All acknowledgments, including those to funding agencies, go at the end of the paper.
Appendix A Appendix
You may include other additional sections here.